OpenStudy (anonymous):

How do I find the horizontal asymptote of 1/(x+7)+3? Show work please.

jimthompson5910 (jim_thompson5910):

\[\Large \frac{1}{x+7}+3\] \[\Large \frac{1}{x+7}+3*\frac{x+7}{x+7}\] \[\Large \frac{1}{x+7}+\frac{3(x+7)}{x+7}\] \[\Large \frac{1}{x+7}+\frac{3x+21}{x+7}\] \[\Large \frac{1+3x+21}{x+7}\] \[\Large \frac{3x+22}{x+7}\] So \[\Large \frac{1}{x+7}+3\] simplifies to \[\Large \frac{3x+22}{x+7}\] Let me know if you can take it from here

OpenStudy (anonymous):

I'm still confused on how to get the horizontal asymptote.

jimthompson5910 (jim_thompson5910):

What are the leading coefficients for \[\Large \frac{3x+22}{x+7}\]

OpenStudy (anonymous):

3 and 1?

jimthompson5910 (jim_thompson5910):

now divide them to get the horizontal asymptote: 3/1 = 3 So the horizontal asymptote is the equation y = 3

OpenStudy (anonymous):

oh ok, thanks so much!

jimthompson5910 (jim_thompson5910):

np